OpenStudy (anonymous):

How do I convert (x-4)^2+(y-3)^2=25 to a polar equation?

OpenStudy (codym251):

Use mathway.com

OpenStudy (codym251):

did you use the website it really helps me with my math question

OpenStudy (anonymous):

Trust me, I did. They're good for everything except this. It really pisses me off.

OpenStudy (john_es):

First thing you should do is expand the equation. Have you try it?

OpenStudy (john_es):

Once you get this, you should try the change: \[x=r\cos\theta \\ y=r\sin\theta\]

OpenStudy (anonymous):

I have, but I only made it as far as \[r^2\cos^2\theta+8rcos \theta+25+r^2\sin^2\theta+6r \sin \theta=25\]

OpenStudy (john_es):

Remember that, in this case, r=5 and that: \[r^2\cos^2\theta+r^2\sin^2\theta=r^2\]

OpenStudy (john_es):

You should simplify what you have correctly obtained.

OpenStudy (anonymous):

\[r^2(\cos^2\theta+\sin^2\theta)=r^2\]

OpenStudy (john_es):

Yes

OpenStudy (john_es):

What is the final result do you get then?

OpenStudy (anonymous):

r^2, I suppose.

OpenStudy (john_es):

No, this is a property you are using to simplify this result you gave me \[r^2\cos^2θ+8r\cosθ+25+r^2\sin^2θ+6r\sinθ=25\]

OpenStudy (john_es):

If you simplify this, you will get this, \[r^2-8r\cos\theta-6r\sin\theta=0\]Ok?

OpenStudy (anonymous):

So this would mean \[r^2-(8rcos \theta - 6rsin \theta)=0\]?

OpenStudy (john_es):

Sorry, your result was wrong with the signs, the correct one is: \[r^2\cos^2θ-8r\cosθ+25+r^2\sin^2θ-6r\sinθ=25\]

OpenStudy (john_es):

Then you get the one I gave, \[r^2−8r\cosθ−6r\sinθ=0\]

OpenStudy (john_es):

Now you can do the following, extract common factor r, \[r(r-8\cos\theta-6\sin\theta)=0\]

OpenStudy (john_es):

So you have two options, \[r=0\] And \[r-8\cos\theta-6\sin\theta=0\] You need the last one, that can be rewritten as, \[r=8\cos\theta+6\sin\theta\] And this is the polar equation.

OpenStudy (john_es):

Do you understand it?

OpenStudy (john_es):

You can check the graph here also, https://www.wolframalpha.com/input/?i=PolarPlot%5B8+cos%5Bx%5D%2B6+sin%5Bx%5D,%7Bx,0,5%7D%5D

OpenStudy (anonymous):

Thanks a million.

OpenStudy (john_es):

You're welcome.